OpenStudy (anonymous):

Write the radical expression 8 over 7th root of x^15 in exponential form. The right answer is 8x^-15/7. Why?

OpenStudy (mathmale):

"8 over 7th root of x^15" comes out as follows:\[\frac{ 8 }{ \sqrt[7]{x ^{15}} }\]

OpenStudy (mathmale):

Focus on the denominator. Rewrite this radical in exponential notation. Here's the rule you need:\[\sqrt[a]{b^c}=b ^{c/a }or. b ^{\frac{ c }{ a }}\]

OpenStudy (anonymous):

Thank you!